Results for this submission Entered Answer Preview Result 7 80 7/80 incorrect The answer above is NOT correct. (1 point) A palindrome is a number (like 535) that is the same if the digits are written in reverse order. If a number is a three digit multiple of eleven then the probability that it is a palindrome is 8/80
Answer #1

There are total 90 three digits palindrome. That is there are 90 three digits palimdrome numbers are between are 100 to 999.

Among these 90 numbers there are, 8 numbers which are multiple of 11:

Those numbers are, 121, 242, 363, 484, 616, 737, 858 and 979

Therefore, the probability that any 3 digits palindrome is also a multiple of 11 is, 8/90 = 4/45 = 0.0889

Answer: 8/90 = 4/45 = 0.0889